OpenStudy (emilyjones284):

I just need help on the solving part of this question... I understand the beginning, ill give a medal(:. Ill post it below.

OpenStudy (emilyjones284):

If \[f(x)=\frac{ 1 }{ 2 }x-6\], what is the equation for \[f ^{-1}(x)\]

hartnn (hartnn):

y = (1/2)x -6 x = (1/2)y-6 add 6 to both sides, what u get ?

OpenStudy (emilyjones284):

x+6=1/2y

hartnn (hartnn):

right, now multiply both sides by 2 ....

OpenStudy (emilyjones284):

2(x+6)... should I distribute the 2 because theres an answer where it is distributed and an answer where its not

hartnn (hartnn):

what do u get after distributing ?

OpenStudy (emilyjones284):

oh wait there is 2(x+6) and 2x+6 so it would have to be 2(x+6) right?

hartnn (hartnn):

yes, after distributing 2x +12 so, 2 (x+6) wouk be correct :)

OpenStudy (emilyjones284):

okay I understand thank you so much (:

hartnn (hartnn):

welcome again ^_^